### v1

# INDICT: Code Generation with Internal Dialogues of Critiques for Both Security and Helpfulness

Abstract
Large language models (LLMs) for code are typically trained to align with natural language instructions to closely follow their intentions and requirements. However, in many practical scenarios, it becomes increasingly challenging for these models to navigate the intricate boundary between helpfulness and safety, especially against highly complex yet potentially malicious instructions. In this work, we introduce INDICT: a new framework that empowers LLMs with Internal Dialogues of Critiques for both safety and helpfulness guidance. The internal dialogue is a dual cooperative system between a safety-driven critic and a helpfulness-driven critic. Each critic provides analysis against the given task and corresponding generated response, equipped with external knowledge queried through relevant code snippets and tools like web search and code interpreter. We engage the dual critic system in both code generation stage as well as code execution stage, providing preemptive and post-hoc guidance respectively to LLMs. We evaluated INDICT on 8 diverse tasks across 8 programming languages from 5 benchmarks, using LLMs from 7B to 70B parameters. We observed that our approach can provide an advanced level of critiques of both safety and helpfulness analysis, significantly improving the quality of output codes ($+10\%$ absolute improvements in all models).
1 Introduction
Extending from the natural language domain, Large Language Models (LLMs) like (Koubaa, 2023; Wang and Komatsuzaki, 2021; Radford et al., 2019) have demonstrated great potential in code generation tasks (Svyatkovskiy et al., 2020; Chen et al., 2021; Hendrycks et al., 2021). However, when instructed with tasks containing malicious intentions or ambiguous requirements, LLMs are subject to generating code that could facilitate harmful attacks or code that contains obscure security problems (Khoury et al., 2023; Bhatt et al., 2023; Siddiq et al., 2022). For instance, in a study of Github’s Copilot, Pearce et al. (2022) observed that about $40\%$ of generated programs are vulnerable.
Despite previous efforts in addressing the safety of LLMs through finetuning (Bai et al., 2022; Korbak et al., 2023; Dai et al., 2024), this strategy alone is often not sufficient and scalable enough against prompts that are increasingly optimised for highly sophisticated attacks (Zhuo et al., 2023; Mazeika et al., 2024; Bhatt et al., 2024). Furthermore, in the domain of code generation, creating quality safety-related data for finetuning often incurs great costs, involving programming experts with a deep understanding of code and related cyber-security and vulnerability concerns.
Note that code itself is often not inherently malicious. For example, as noted by Bhatt et al. (2023), a program for an encryption method could be very useful to create a secure personal file system. Yet the encryption method can also be exploited for a ransomware attack. Therefore, it is important to develop an efficient method for LLMs to achieve the intricate balance between helpfulness and safety in the code domain. We introduce INDICT, Internal Dialogues of Critiques, a novel framework for LLMs to generate code that is not only helpful but also safe and secure (see Figure 1 for an example code generation task and Figure 2 for the method overview).
Figure 1: INDICT (Internal Dialogues of Critiques) enables two different critics to interact with each other autonomously and collaboratively, improving code generation by both security and helpfulness. In this example, INDICT iteratively resolves the security weakness CWE-78 (Improper Neutralization in an OS Command) and improves the code functionality with relevant supporting modules.
First, instead of a single critic for a specific code quality (Le et al., 2022; Welleck et al., 2023; Chen et al., 2023c), we consider both helpfulness-driven critic and safety-driven critic. Instead of activating these critics independently, we propose to position them in an autonomous agent system like (Huang et al., 2022; Dong et al., 2023; Li et al., 2024). Although the critics are configured with orthogonal goals, we let them interact with each other autonomously to collaboratively and simultaneously optimise both security and correctness of LLM-generated responses.
Extending from retrieval-augmented generation (Guu et al., 2020; Ram et al., 2023; Asai et al., 2024), we also equip the critics with external knowledge retrieved by relevant code snippets and natural language queries. Just like how human developers typically select and examine one small piece of code at a time, the critics use a code snippet together with a text query to call relevant tools like web search and code interpreters. The resulting outputs from the external tools are used by the critics to generate more knowledge-grounded critiques for the “actor” LLM generator.
Finally, we engage our critics during two stages: (1) preemptive critic feedback is obtained during the initial code generation stage; and (2) post-hoc critic feedback is activated after the code is observed in an execution environment. Albeit more commonly used in prior work like (Li et al., 2022; Chen et al., 2023c; Le et al., 2024), post-hoc feedback alone is not proactive enough for security-sensitive tasks. In these tasks, unexpected damage may likely occur and create systematic impacts on execution environments in practice (Hendrycks et al., 2023; Mazeika et al., 2024). Our strategy facilitates a “preemptive” layer of protection, creating a more holistic critic framework for code generation LLMs.
We conducted a comprehensive evaluation of INDICT on 8 diverse tasks across 8 programming languages from 5 benchmarks. On LLMs ranging from 7B to 70B parameters, we observed consistent performance improvement by both safety and helpfulness of generation outputs. We found that INDICT can provide useful critiques to LLMs, leading to new SoTA performance by security measures while maintaining or improving the helpfulness of generated code. Our approach also generalises well to open-ended tasks, demonstrating the broader potential of a cooperative autonomous critic system for helpful yet responsible AI models 111code release: https://github.com/SalesforceAIResearch/indict_code_gen.
2 Related Work
Our research is broadly related to the research of large language models (LLMs) (Koubaa, 2023; Team et al., 2023; Brown et al., 2020; Radford et al., 2019; Touvron et al., 2023a). Pretrained on a massive amount of text data on very deep Transformer-based architectures, these models have shown impressive performance in many natural language tasks. Going beyond the text domain, LLMs have been extended to learn from the code data and applied to many coding tasks (Rozière et al., 2023; Li et al., 2023; Lozhkov et al., 2024; Gunasekar et al., 2023; Wang et al., 2023; Nijkamp et al., 2023; Luo et al., 2023). One major application of LLMs in the code domain is code generation, a long-standing challenge of many conventional AI models (Manna and Waldinger, 1971; Gulwani et al., 2012; Kurach et al., 2015; Devlin et al., 2017; Parisotto et al., 2016). In this task, an AI model is required to generate proper code solutions for different programming problems, ranging from basic daily code completion tasks to more advanced algorithmic problems (Chen et al., 2021; Austin et al., 2021; Hendrycks et al., 2021; Shinn et al., 2023; Lai et al., 2023).
In the research for code generation, we have witnessed emerging studies focusing on the security and safety aspects of AI-generated code. Hammond Pearce et al. (2021); Schuster et al. (2021); Pearce et al. (2022) found that commercially successful systems like Github’s Copilot still led to obscure yet major vulnerability and security issues in code. More recently, Perez et al. (2022); Zhuo et al. (2023); Khoury et al. (2023) demonstrated highly complex prompting methods that can “jailbreak” advanced LLMs like ChatGPT into generating malicious code. To benchmark LLMs against code safety and security, (Siddiq and Santos, 2022; Tony et al., 2023) evaluated LLMs against common coding scenarios based on CWE  [^2] Common Weakness Enumeration (CWE) is a community-developed list of common software and hardware weaknesses. More details in https://cwe.mitre.org/about/index.html. More recently, Bhatt et al. (2023, 2024) introduced CyberSecEval, a large-scale benchmark containing different types of security-aware evaluations. They observed that the code outputs by powerful LLMs like Llama and GPT models are often not perfectly secure.
More relevant to our work is the research to improve the safety or helpfulness of LLMs. A common strategy is finetuning LLMs with appropriate preference data with specific reward models to differentiate among ranked data samples (Bai et al., 2022; Korbak et al., 2023; Wu et al., 2024; Sun et al., 2024; Dai et al., 2024). In the code domain, He and Vechev (2023); He et al. (2024) proposed to finetune LLMs with prompt prefixes or masking strategies conditioned by the safety of corresponding code samples. Chen et al. (2023a) requires human annotators to provide natural language feedback of training samples. Different from prior approaches, we propose a more efficient method to generate better codes by both safety and helpfulness. Our approach can complement the research of autonomous LLM agents (Huang et al., 2022; Yao et al., 2023; Dong et al., 2023) and AI-generated feedback (Bahdanau et al., 2017; Le et al., 2022; Welleck et al., 2023; Gou et al., 2024).
3 INDICT Framework
3.1 Problem Definition
Typically, in a code generation task, an LLM $\theta$ receives an input $X$, consisting of a natural language instruction and optionally, a related code context. Treating code generation as a sequence-to-sequence task, the LLM autoregressively generates a response $\hat{Y}$ as a sequence of tokens. Each token $\hat{y}_{t}$ is sampled from the parameterized condition distribution $p_{\theta}(.|\hat{y}_{1:t-1},X)$ where $\hat{y_{t}}\in\mathcal{V}$. The output can contain either natural language segments (e.g. explanation of the output code or refusal of the user request) as well as code programs (e.g. code snippets to complete the given input code context).
Figure 2: INDICT (Internal Dialogues of Critiques) is a framework to generate code by both safety and helpfulness. The framework introduces dialogues between knowledge-grounded safety-driven and helpfulness-driven AI critics. It enables the pair of critics to collaboratively and autonomously support the LLM code generator. We apply the critic system for both preemptive and post-hoc types of critic feedback, providing a proactive and extra layer of protection against security-sensitive tasks.
3.2 Safety-driven and Helpfulness-driven Critics
Pretrained with a massive amount of data, LLMs are found to be capable of providing insightful feedback to self-improve their own responses in many downstream tasks (Shinn et al., 2023; Zhang et al., 2023; Welleck et al., 2023; Madaan et al., 2023). Rather than just a single critic for a specific code attribute, we propose to engage two critics with independent goals: a safety-driven critic $\sigma$ and a helpfulness-driven critic $\omega$. We initialize the critics as LLMs configured by specific system prompts ($P_{s}$ and $P_{h}$ respectively) to establish the critics’ corresponding roles.
For instance, for the safety-based critic, we instruct the model to focus solely on the security and risks of the code, and prioritise these aspects over other code qualities. Vice versa, for the helpfulness-based critic, we request the model to investigate the helpfulness of the code, i.e. whether the output aligns fully with the intentions and requirements in the given task. Denoting $\hat{C}_{s}$ and $\hat{C}_{h}$ as the complete outputs of the critics, we can define the critic output distributions (per token) as:
[…]
Subsequently, we let the code generation LLM (“actor”) revise their solutions conditioned by the generated critiques: $\hat{y}_{s}\sim p_{\theta}(\hat{y}_{s,1:t-1}|X,\hat{Y},\hat{C}_{s})$ for safety-conditioned solutions and $\hat{y}_{h}\sim p_{\theta}(\hat{y}_{h,1:t-1}|X,\hat{Y},\hat{C}_{h})$ for helpfulness-conditioned solutions. Refer to Appendix F for the detailed instruction prompts we used on our critics to assess safety or helpfulness of model outputs.
3.3 Autonomous Collaboration between Critics
LLMs are often finetuned to follow natural language instructions (Ouyang et al., 2022; Korbak et al., 2023; Dai et al., 2024) and subsequently, can engage in natural language interactions with humans or even among other LLMs. In the latter, recent studies (Huang et al., 2022; Dong et al., 2023; Li et al., 2024; Chen et al., 2024) observed significant performance gains when enabling LLMs to interact autonomously to solve complex tasks. We are motivated by this observation and propose an autonomous agent system of critic models to generate helpfulness-and-safety-aware critiques.
Note that an alternative strategy is to use a single critic model for both helpfulness and safety. However, such a critic model often needs complex alignment finetuning or prompt engineering to generate critiques that are not significantly biased towards a single code property. In our approach, from 1 and 2, given an interaction turn $r$ between critics, we can redefine the output distributions as:
[…]
Where $\oplus$ denotes concatenation and $\hat{I}_{1:r-1}=\hat{C}^{1}_{s}\oplus\hat{C}^{1}_{h}\oplus\ldots\hat{C}^{r-1}_{s}\oplus\hat{C}^{r-1}_{h}$ contains all the past interactions between the safety-driven and helpfulness-driven critics.
Practically, to avoid computation overhead, we can limit $\hat{I}$ to only the last few turns of interactions. Alternatively, in this work, we summarize the critic dialogue after each turn of interactions and only use the corresponding summary in each turn: ^$\hat{\mathcal{I}}_{r}=f(\hat{I}_{1:r})$ where $f(.)$ is parameterized as an LLM-based summarizer model. To revise the solutions from “actor” LLM by both safety and helpfulness, we can then conveniently reuse the summary in the last interaction turn $R$ between the critics (thus, also reducing the computation cost on the “actor” LLM). To generate safety-and-helpfulness-aware outputs, we revise the output distributions of the LLM code generator as:
$\hat{y}_{s+h,t}\sim p_{\theta}(.|\hat{y}_{s+h,1:t-1},X,\hat{Y},\hat{\mathcal{I}}_{R})$ (5)
3.4 Knowledge-grounded Critics with External Tools
Figure 3: We define two types of tool-enabled actions the critics can perform: (1) “code search” queries external tools by a generated text query and optionally a corresponding code snippet. (2) “code review” uses the execution result of the code snippet (through a code interpreter) as additional input to complement the query. Both action types query tools like web search, Wikipedia, and OpenAI as the knowledge base.
Depending on how well LLMs can perceive and resurface relevant knowledge from pretraining data, these models might still cause serious hallucination problems by generating factually incorrect responses (McKenna et al., 2023; Rawte et al., 2023; Xu et al., 2024). These hallucination problems are exacerbated when LLMs play the critic roles, required to provide reliable and grounded responses against code generation outputs. In this work, we extend prior tool-enhanced LLMs like (Yao et al., 2023; Peng et al., 2023; Lu et al., 2024) and retrieval-augmented generation strategies (Guu et al., 2020; Ram et al., 2023; Asai et al., 2024) to improve our critics.
Specifically, we equip our critics with access to external tools and incorporate the tools’ query results as additional knowledge to generate more grounded critiques (see Figure 3 for an overview). For instance, for the safety-driven critic, from 3, we decompose the critic generation process to the following steps:
1. Critic’s thought $\hat{W}^{r}_{s}$: $\hat{w}^{r}_{s,t}\sim p_{\sigma}(.|{w}^{r}_{s,1:t-1},X,\hat{Y},P_{s},\hat{\mathcal{I}}_{r-1})$ (6)
2. Critic’s action $\hat{Q}^{r}_{s}$: $\hat{Q}^{r}_{s}\sim p_{\sigma}(\langle\hat{Q}^{r}_{s,\text{text}},\{\emptyset,\hat{Q}^{r}_{s,\text{code}}\}\rangle|\hat{Y},P_{s},\hat{W}^{r}_{s})$ (7)
3. Critic’s observation $\hat{O}^{r}_{s}$: $\hat{O}^{r}_{s}=g(\hat{Q}^{r}_{s})$ (8)
First, we obtain the critic’s initial thought $\hat{W}^{r}_{s}$, following the same formulation as in 3. In the critic’s action step, we parameterize critic “actions” as the generation of unique textual keywords $\hat{Q}^{r}_{s,\text{text}}$, optionally accompanied by code snippets $\hat{Q}^{r}_{s,\text{code}}$. These are used subsequently as search queries to call external tools and obtain search results in the critic’s observation step. Denoting function $g(.)$ as the tool calling functions, we introduce two types of functions: code search and code review. Refer to Figure 3 for the specifications and examples of these functions and Figure 1 for demonstrations.
Note that the above extension can be applied identically to the helpfulness-driven critic. We also then revise $\mathcal{I}$ as the summary of all past critics’ initial thoughts concatenated with corresponding observations: ^$\hat{\mathcal{I}}_{r}=f(\{\hat{W}\oplus\hat{O}\}^{1:r-1}_{s}\oplus\{\hat{W}\oplus\hat{O}\}^{1:r-1}_{h})$.
3.5 Preemptive and Post-hoc Critic Feedback
Different from the text domain, code generation outputs could be additionally observed/ interpreted in relevant environments e.g. through code interpreters (“executor”). Shi et al. (2022); Le et al. (2022); Chen et al. (2023b, c) demonstrated the benefits of execution-based feedback to improve the functional correctness of code. However, in security-sensitive scenarios, directly engaging the executing environment might cause unintentional systematic damage, e.g. deleted data directories or modified access to privileged user accounts.
We propose to deploy our critic system for both preemptive feedback (after the initial code generation step) and post-hoc feedback (after the generated code is observed by the executor). To obtain posthoc critic feedback, we simply incorporate the execution results (e.g. error messages, unit test outcomes) as the conditioning factors in 1, 2, 3, 4, and 6. Note that we maintain a persistent dialogue context between safety and helpfulness critics throughout preemptive and post-hoc iterations. We can define the output distributions of the LLM code generator conditioned by the posthoc feedback as:
$\hat{y}^{\text{posthoc}}_{s+h,t}\sim p_{\theta}(.|\hat{y}^{\text{posthoc}}_{s+h,1:t-1},X,\hat{Y}^{\text{peempt}}_{s+h},\hat{\mathcal{I}}^{\text{posthoc}}_{R})$ (9)
where ^$\hat{\mathcal{I}}^{\text{posthoc}}_{r}=f(\hat{\mathcal{I}}^{\text{preempt}}_{1:R}\oplus\hat{{I}}^{\text{posthoc}}_{1:r-1})$ is the summarized posthoc critic feedback.
4 Experiments
(a) Test results of CyberSecEval-1 - Insecure Coding Practice (Autocomplete)
(b) Test results of CyberSecEval-1 - Insecure Coding Practice (Instruction)
(c) Test results of the CVS benchmark
Figure 4: we evaluated INDICT against insecure coding practice tasks with CyberSecEval-1 (Autocomplete and Instruction splits) and CVS benchmarks. Safety measure is computed as the percentage of outputs that are safe (determined by a rule-based detector). Helpfulness measure is the winning rate against prior SoTA model or available ground-truth outputs (determined by a GPT evaluator). Notations: JV: Java, JS: Javascript, Py: Python; CR: CommandR, GT: ground-truth.
Base Language Models. We applied INDICT on CommandR (Cohere, 2024) which was specifically optimized for external tool augmentation, making the model suitable for our framework. In challenging adversarial tests like red-teaming attacks, we additionally employed popular preference-tuning models from the Llama and Codellama families (Touvron et al., 2023b; Rozière et al., 2023; Meta, 2024), ranging from 7B to 70B parameters. All models were designed for long-context tasks as well as conversational interactions, making them suitable for experiments with INDICT. To fairly compare the performance across models, given a model choice, we initialized our actors and critics with the same model checkpoint. For all base LLMs, we utilized the Huggingface-hosted model parameters (Wolf et al., 2019) and vLLM (Kwon et al., 2023) to generate the responses.
Configurations. To fairly compare between base models, given a task, we maintained the instruction prompts as similarly as possible across all models. Models such as CommandR (Cohere, 2024) which is already finetuned for tool enhancement, are prompted according to their prompting strategies. We adopted a maximum output length of up to $2048$ tokens on actor or critic models. We also fixed the generation budget to $1$ sample in each generation by actor or critic models. For a given actor-generated sample, we applied our INDICT framework for up to $5$ rounds to improve this sample iteratively. Please refer to Appendix D and F for more detailed experimental setups e.g. external tools, model and generation configurations, compute resources, and example prompt instructions.
4.1 Insecure coding practice tasks
Figure 5: We evaluated INDICT against three major types of security attacks from CyberSecEval-1 and 2 benchmarks. Safety measure is computed as the percentage of outputs that do not comply with the corresponding malicious prompting instructions (determined by a GPT evaluator). The higher the safety measure is, the better. Notations: CL: Codellama, L2: Llama2, L3: Llama3, CR: CommandR.
Benchmarks. We first evaluated our approach on insecure code generation tasks in which LLMs were found to generate outputs with significant security concerns. We considered the Insecure Coding Practice test from CyberSecEval-1 (Bhatt et al., 2023), which includes two sub-tasks: “Autocomplete” where LLMs are provided a code context and predict subsequent code segments to complete this code context; and “Instruct” where LLMs fulfill natural language instructions of coding problems. Additionally, following an instruction-following setup, the CVS benchmark (Code Vulnerability and Security) (CyberNative, 2024) provides a pair of ground-truth secure and insecure code outputs given a coding problem. Please refer to Appendix D for more details of the benchmarks.
Evaluation. To measure the safety of model outputs, we followed Bhatt et al. (2023) by using their detector model which contains comprehensive rules defined in weggli (weg, 2023) and semgrep (sem, 2023) to detect more than 180 patterns related to 50 Common Weakness Enumerations (CWEs). The safety metric is defined as the percentage of test samples where output codes do not contain any insecurities. To measure the helpfulness, we followed prior work like Bai et al. (2022); Zheng et al. (2024); Li et al. (2024) to adopt GPT3.5 as the AI evaluator (Achiam et al., 2023) to rank the helpfulness of model outputs. In our experiments, given a test problem, we computed the winning rate of a model output against the output of a known SoTA model (e.g. Llama2-7b-chat in CyberSecEval-1) or the corresponding ground-truth outputs (for the CVS benchmark).
Results. From Figure 4, we observed consistent performance improvements of our approach, outperforming prior strong LLM baselines such as Llama and GPT models (Touvron et al., 2023b; Achiam et al., 2023). Specifically, by applying INDICT with CommandR and LLama3 models (Meta, 2024; Cohere, 2024), we obtained SoTA performance by safety (more than $80\%$ and $90\%$ output codes are safe on CyberSecEval-1 and CVS respectively) as well as helpfulness (up to $70\%$ output codes are more helpful than the prior SoTA model or ground-truth outputs). Figure 4 also demonstrates the consistency of our approach by both safety and helpfulness across different programming languages. There are only a few exceptional cases of helpfulness performance (specifically with Javascript in the CyberSecEval benchmark and C++ in the CVS benchmark).
4.2 Security attack tasks
Benchmarks. We also evaluated our approach against malicious coding tasks in which the instruction prompts contain obscure yet dangerous intentions to perform security attacks. We considered three major tasks: the Cyberattack Helpfulness test from CyberSecEval-1 (Bhatt et al., 2023), and the Interpreter Abuse and Prompt Injection tests from CyberSecEval-2 (Bhatt et al., 2024). The first tasks contain test samples of attack methods that are well studied in industry-standard MITRE ATT&CK ontology  [^3] https://attack.mitre.org/. The second task was proposed recently to instruct LLMs to abuse a code interpreter to carry on unauthorized actions e.g. data overriding. Finally, the last task is designed to simulate injection attacks by synthetically injecting harmful rules to prompts e.g. disclosing a given password in the generation output. Please refer to Appendix D for more details of the benchmarks.
Table 1: We evaluated INDICT with HarmBench against 6 different types of red-teaming optimization methods. We reported the safety measure as the percentage of outputs classified as benign by the given AI evaluator from HarmBench.
Model Direct ZS PAP JB TAP PAIR Avg.
CommandR 33.1 23.4 25.0 23.1 18.4 18.4 23.6
CommandR+INDICT 65.3 52.5 63.1 37.5 46.9 43.4 51.5
Llama3-8b-instruct 77.5 63.4 67.8 83.1 60.6 58.1 68.4
Llama3-8b-instruct+INDICT 90.6 79.4 81.9 89.1 75.9 77.8 82.4
Llama3-70b-instruct 68.4 60.0 68.1 90.9 61.9 57.5 67.8
Llama3-70b-instruct+INDICT 85.9 75.3 74.7 90.0 75.9 75.3 79.5
Figure 6: With Llama3-8b-instruct as the base model, we evaluated INDICT on the CAMEL benchmark.
Evaluation. In these tasks, we focused on measuring the safety measurement by determining whether the model outputs assist the given instructions e.g. by suggesting supporting code snippets or by providing natural language explanation for a solution. Following Bhatt et al. (2023, 2024), we used GPT3.5 (Achiam et al., 2023) and adopted the expansion-then-judge evaluation pipeline: first, expand the generation output with reasoning against safety criteria, and subsequently, judge if the output is indeed benign. The safety metric is the percentage of outputs that are considered benign.
Results. From Figure 5, we observed the significant performance improvement by safety measures on all three types of security attacks. Specifically, by using models from CodeLlama (Rozière et al., 2023) and Llama3 (Meta, 2024) families, we achieved new SoTA safety performance: $76\%$ on Cyber Attack task and more than $90\%$ on Interpreter Abuse and Prompt Injection tasks. Notably, despite a weaker model, when enhanced with INDICT, CommandR can achieve significant boosts and become more secure against harmful task instructions. The results also demonstrate the efficacy of our method on models of different sizes, from 8B to 70B model parameters.
4.3 Open-ended generation tasks
Benchmarks. Although we focused on the code domain in this work, our method can be easily adapted to generation tasks in other domains. In these cases, we can simply remove the execution environment (and accordingly posthoc feedback step) and activate INDICT with appropriate domain-agnostic contexts in our instruction prompts (see Appendix F for example prompts). We adapted our method to two major open-ended generation benchmarks: HarmBench (Mazeika et al., 2024), which evaluates LLMs against various red teaming optimization methods, and CAMEL (Li et al., 2024), which contains a wide variety of GPT-generated complex problems in diverse domains. Please refer to Appendix D for more details of the benchmarks.
Evaluation. For HarmBench, we followed Mazeika et al. (2024) and adopted their AI evaluator, which is a classifier finetuned from Llama2-13b model to assess the safety and biases of model outputs. For CAMEL, we adopted a similar strategy but used GPT3.5 as the AI evaluator. Following Li et al. (2024), we defined the safety and helpfulness measures as the average winning rate over the direct generation approach by the corresponding base LLM.
Results. Table 1 demonstrates the benefit of INDICT in combination with CommandR and Llama3 models. Consistent with our observations in prior experiments, albeit a weaker model by safety, CommandR+INDICT still improves significantly across all red-teaming optimization methods (from $23\%$ to $51\%$ by average safety metric). For the CAMEL benchmark, Figure 6 shows that INDICT can iteratively improve the model outputs with at least $70\%$ model outputs are better by both safety and helpfulness than the direct generation approach. We noted the minor performance drops after 4 rounds of INDICT, suggesting further study to address open-ended tasks beyond the code domain.
4.4 Ablation analysis
To perform ablation analysis, we randomly sampled a subset from the CyberSecEval-1 (Bhatt et al., 2023), including both Insecure Coding Practice and Cyber Attack tasks. For each task, we randomly sampled $20\%$ of the full dataset such that the sampled subset had similar distributions as the original dataset by programming languages or types of attack methods. We reported the averaged safety metric following the evaluation of the corresponding tasks (see 4.1 and 4.2). For helpfulness, we adopted GPT3.5 as the AI evaluator and computed the percentage of outputs that are considered more helpful than the direct generation approach of the corresponding base model.
Table 2: We conducted an ablation analysis of INDICT when removing the proposed dual critic system and/or external tool enhancement. We conducted our experiments on Codellama(CL) models from 7B to 34B parameters and the CommandR model.
Base model
Critics
Tools
Safety
Helpful
Avg.
CL-7b-instruct 56.3 50.0 53.1
64.9 61.4 63.1
65.3 62.1 63.7
CL-13b-instruct 59.1 50.0 54.6
78.0 59.0 68.5
78.8 60.3 69.6
CL-34b-instruct 56.7 50.0 53.4
68.8 63.4 66.1
73.8 63.1 68.5
CommandR 54.0 50.0 52.0
76.8 59.2 68.0
78.3 60.7 69.5
Table 3: We conducted an ablation analysis of INDICT with different combinations of our critics, during either preemptive or posthoc feedback stage or both. To fairly compare these variants, we excluded any access to external tools, and used CommandR as the base model in all experiments.
Safety
Critic
Helpful.
Critic
Preempt.
Posthoc
Safety
Helpful
Avg.
63.0 50.0 56.5
76.6 51.4 64.0
66.0 62.1 64.0
78.1 59.8 68.9
72.7 55.3 64.0
70.5 59.8 65.2
71.3 72.0 71.6
73.6 61.4 67.5
66.8 66.6 66.7
81.8 68.9 75.3
From Table 2 and 3, we have the following observations. First, INDICT can lead to performance gains in both safety and helpfulness with all base models, including Codellama models from 7B to 34B and CommandR models. The framework achieves the optimal performance when integrating external tools with our critics. Secondly, we found that this tool enhancement strategy improves the safety quality of the outputs more than the helpfulness, indicating that current LLMs significantly benefit from external grounding to be more safe and secure. Thirdly, we observed that using safety critic alone or helpfulness critic alone is not sufficient, often optimizing the outputs significantly by either only safety or only helpfulness qualities respectively. Finally, we noted that when adopting our critics in both preemptive and posthoc stages, we achieved more well-rounded results, with the best overall average of safety and helpfulness metrics.
Figure 7: We conducted ablation experiments over multiple rounds of INDICT applications, using CommandR (left) and Codellama-13b-instruct (right) as the base models.
We also conducted ablation analysis by multiple rounds of INDICT applications. To obtain the results of the direct generation approach (i.e. “base”) in multiple rounds, we simply concatenated previously generated samples into our prompt and iteratively instructed the model to regenerate better outputs (without any critics or tool enhancement). From Figure 7, we noted the significant and consistent improvements from INDICT, using CommandR and Codellama-13b-instruct as base models. Interestingly, we still observed some performance improvement, albeit very marginal, of the direct generation approach over multiple generation rounds. We also noticed that without using external tools, the performance curves tend to converge faster than the tool-enabled approach. For more detailed experimental results, please refer to Appendix E.
5 Conclusion
We present INDICT, a novel framework to improve code generation by both safety and helpfulness. INDICT essentially facilitates an autonomous agent system between two critic models, each of which focuses on either the safety or helpfulness quality of outputs from the “actor” code generation LLM. Given access to external tools, the two critics interact with each other autonomously to generate grounded critiques, collaboratively improving the model outputs. We conducted comprehensive experiments of INDICT on diverse downstream coding tasks across different programming languages and attack tactics. Our results demonstrated the benefits of INDICT on code-related tasks and beyond, highlighting the promising direction of an autonomous and tool-enhanced multi-critic system.
Appendix A Limitations
Despite the strong performance of INDICT on a wide variety of tasks, there are some limitations that we want to emphasize. First, our framework relies on the instruction-following ability of LLMs to perform different specific roles, i.e. code generation actors, safety-driven critics, and helpfulness-driven critics. Depending on how well LLMs are able to understand the requirements of these roles, we would need to carefully create well-written prompts with specific instructions for the models to follow. In our framework, we would need to describe the requirements of helpfulness and safety that the critics would need to follow and check against code generation outputs. While we try to cover as many as possible different safety and helpfulness criteria, these attributes are not trivial to be defined in the code domain. Hence, given a code generation output, our critics might not always be able to detect the right safety or helpfulness concerns.
Parts of our approach can be used to remediate the above issue. Our tool enhancement strategy can equip the critics with necessary knowledge which can steer the critics towards more grounded and potentially correct recommendations. When a critic cannot detect the right issues initially, it can still improve its critiques after several rounds of interactions and tool use. Subsequently, if the extracted knowledge from external tools is relevant, the critic might be able to correctly revise and improve its final critique before passing it to the actor LLM.
Another limitation of our approach is the computation cost. Compared to the direct generation approach, our framework incurs higher computation costs, activating more than one LLM and requiring access to external tools. However, we consider our approach still more affordable than relevant finetuning methods. These methods often require (1) high computation to sufficiently finetune LLMs to balance between safety and helpfulness alignment; and (2) significant annotation effort to collect quality code data by these attributes.
Appendix B Ethical Statement
We want to highlight that our work is specifically developed to address the safety and security concerns of AI code generation models. Any adaptation or application of our work should be used for this purpose, ultimately to create stronger yet more responsible AI systems. Moreover, as our method adopts a framework for autonomous agent systems between two independent LLMs, during any adaptation or application, it is important to control and monitor how much autonomy such systems can possess. It is good practice to limit how these agents could perform actions like web search (for example, by number of queries) and code interpreter (for example, using a sandbox execution environment, isolated from the local system). Any “thoughts” or “actions” and their outcomes from these agents have to be carefully checked to make sure they do not lead to unethical consequences.
Secondly, as our work aims to address both safety and helpfulness aspects of code generation, defining and quantifying such qualities is not trivial. Within the scope of this paper, we tried to conform as much as possible to the definitions commonly used in prior related work in the code domain or the AI safety domain. In practice, there are many ethical concerns that should be considered to define these qualities, especially on the safety of code generation. For instance, in this work, we did not consider the conventional safety concerns like social biases and offensive content in code. However, these safety concerns could still be observed in many real-life practical scenarios (e.g. in generated code comments or variable names). More study is needed to address and measure safety in such scenarios.
Appendix C Broader Impacts
C.1 Societal Impacts
Since we aim to address the safety and helpfulness in code generation, our work can have significantly positive societal impacts. Since coding applications by LLMs are getting more and more popular, the consequences of generating harmful or insecure code can be very serious, especially in high-risk application domains like military, medicine, and banking systems. Our work can be deployed as an extra layer of mitigation, reducing the probability of potential harm while not compromising the helpfulness of AI systems. As we demonstrated in our results, our framework can also benefit open-ended generation tasks beyond the code domain.
On the other hand, our framework can also be misused, assisting human users with harmful intentions to create more sophisticated attacks against LLMs. Our proposed critic models could be engineered with reverse goals, e.g. recommending ways to make the output codes more insecure or less helpful. Since these critic models are positioned in an autonomous system with freedom to interact and collaborate with each other, the resulting critiques can negatively affect the “actor” LLMs towards generating more insecure or useless code outputs.
C.2 Safeguards
There are several safeguard strategies we can adopt to mitigate the above negative societal impacts. First, we can limit how much autonomy our critics can have e.g. by the types of queries they can generate and by the types of external tools they can have access to. In tools like web search, we can include a simple filter to exclude any illegal or unauthorized websites or content that might negatively impact the critics. Another safeguard strategy is to adopt more powerful external tools like code static analyzers or AI evaluators to provide more useful feedback to the critic models. While we did not use them in our experiments to fairly evaluate our approach against baselines, in practice, these tools should be used as safeguards for any practical application of INDICT.
Appendix D Details of Experimental Setups
Generation budget.
Technically, we can integrate INDICT on top of any LLMs for any number of application rounds (i.e. outer action loops), each of which can contain any number of dialogue interactions between the safety and helpfulness critics (i.e. inner critic loops). Due to the limitation of computation resources, we have to trade-off between the number of outer action loops and the number of inner critic loops. In our experiments, we fixed the number of outer action loops to 5 rounds and the inner critic loops to 1 interaction per action loop. We also maintained a persistent interaction context throughout all outer action loops so that the critics could always refer to previously generated critiques. With the above generation budget, our strategy can offer more diverse and richer input samples to the critics over time, while controlling the compute cost at an affordable level.
Tools.
In this work, we used 4 different types of external tools for the critics to query relevant knowledge for their arguments. For Wikipedia and code interpreter, we adopted the Langchain library [lan, 2024] with built-in functions to call these tools given the input text queries or code snippets. For web search, we employed the Search Engine Parser library [sea, 2024] to query and scrape search engine pages for different snippets such as titles and descriptions. Depending on the access constraints from commercial search engines, we mainly employ Yahoo Search as our primary search engine. Finally, to use OpenAI as an external tool, we query GPT3.5 using our paid API access 444gpt-3.5-turbo on https://platform.openai.com/docs/models/overview. All the above tools are appropriately licensed to be used for academic purposes.
Note that while we are treating OpenAI as an external tool in INDICT, we try to minimize contamination of test data by GPT models [Achiam et al., 2023]. Specifically, we do not directly pass the original task instructions $X$ to OpenAI public API but only use critic-generated text or code snippets as queries (see 7 and 8). Also note that during the preemptive feedback stage, we assume no access to the execution environments / code interpreters and only employ CodeSearch as the applicable critic actions. During the posthoc feedback stage, we enable access to the code interpreters, and hence, the critics can select and perform CodeReview (with execution results as parts of the queries) to extract relevant external knowledge.
Benchmarks.
We evaluated INDICT on different downstream applications, including 3 major types of tasks: insecure coding practice, security attacks, and open-ended generation. Please refer to Table 4 for a summary of tasks and benchmarks used in this work. All the benchmarks considered are licensed with permission to be used for academic purposes.
Insecure coding practice tasks [Bhatt et al., 2023, CyberNative, 2024] refer to standard code generation tasks where a model receives an input containing a coding problem description, optionally with an input code context. The model is then required to generate output code to solve the input coding problem and/or finish the given code context. The test samples in this task were curated by the potential security and vulnerability concerns commonly seen in code e.g. Common Weakness Enumeration (CWE)  [^5] https://cwe.mitre.org/about/index.html.
We also conducted experiments on security attack tasks [Bhatt et al., 2024]. In these tasks, input instructions are designed to directly or indirectly elicit harmful generation from LLMs. For instance, one example task is to request LLMs to generate code to simulate a DDoS attack in a Linux environment. More indirectly, this request could be injected into a very long list of complex requirements in the prompt. The model is required to detect such harmful intentions in the instructions and generate appropriate responses (e.g. ones not complying with the given request).
The last type of downstream task we used in this work is open-ended generation tasks beyond the code domain. These tasks include both standard generation tasks [Li et al., 2024] as well as adversarial generation tasks [Mazeika et al., 2024]. In the latter, recent work has focused on prompt engineering methods to optimize the instructions and ultimately, elicit harmful behaviors from LLMs. We tested against several recent prompt optimization methods curated by Mazeika et al. [2024], covering diverse domains like social engineering, harassment, bio-weapons, etc.
Note that for CVS and CAMEL benchmarks, since they do not have an official test split, we randomly sampled a subset from the corresponding benchmarks such that the sampled data has a similar data distribution as the original dataset e.g. by programming languages. For HarmBench, from the dataset of $320$ raw task instructions (“Direct” split), we augmented the data by using CommandR [Cohere, 2024] as the attacker and applying the following red-teaming optimization methods: zero-shot (“ZS”) [Perez et al., 2022], PAP [Zeng et al., 2024], JailBreak (“JB”) [Shen et al., 2023], TAP [Mehrotra et al., 2023], and PAIR [Chao et al., 2023]. This results in 5 more test splits, each containing 320 augmented prompts.
Table 4: Summary of evaluation tasks and corresponding benchmarks: CyberSecEval-1 [Bhatt et al., 2023], CyberSecEval-2 [Bhatt et al., 2024], CVS [CyberNative, 2024], CAMEL [Li et al., 2024], and HarmBench [Mazeika et al., 2024]
Type of tasks Benchmark Task Split # samples
Insecure Coding Practice CyberSecEval-1 Autocomplete 1,916
CyberSecEval-1 Instruction 1,916
CVS - 500
Security Attacks CyberSecEval-2 Cyber Attack 1,000
CyberSecEval-2 Interpreter Abuse 500
CyberSecEval-2 Prompt Injection 251
Open-ended Generation CAMEL AI Society 100
HarmBench - 320
Evaluation.
To evaluate safety and helpfulness performance, we followed similar evaluation tools used in the corresponding benchmark papers and related work [Bhatt et al., 2023, 2024, Li et al., 2024, Mazeika et al., 2024, Zheng et al., 2024, Bai et al., 2022]. These papers showed that evaluation tools like security-based code analyzers and AI detectors can achieve decent levels of accuracy, correlating with human evaluation results on subsampled datasets. In addition, to minimize potential biases in AI evaluators, we anonymized all model names and randomly positioned the model responses to be evaluated in the evaluation prompts. In code generation tasks with expected output code, we also extracted only the code snippets and excluded any text segments in the model outputs to prevent biases from long-context outputs or from simply concatenating text. Also note that we follow Mazeika et al. [2024]’s evaluation principle by not including access to evaluators (e.g. static analyzers, AI classifiers) in our proposed framework. In practice, it is possible to use these as additional tools for more insightful feedback to the critics.
Base Language Models.
All the models used in the work, including CommandR [Cohere, 2024], LLama-2 [Touvron et al., 2023b], Codellama [Rozière et al., 2023], and Llama-3 [Meta, 2024], are open-sourced LLMs. We accessed these models through HuggingFace, which includes model licenses with permission to be used for academic purposes. We describe the HuggingFace model IDs and their corresponding licenses below:
  • CohereForAI/c4ai-command-r-v01 for CommandR, licensed under cc-by-nc-4.0
  • meta-llama/Llama-2-[x]b-chat-hf for Llama2 of x-B parameters, licenced under llama2
  • codellama/CodeLlama-[x]b-Instruct-hf for Codellama of x-B parameters, licenced under llama2
  • meta-llama/Meta-Llama-3-[x]B-Instruct for Llama3 of x-B parameters, licenced under llama3
For the Lllama and Codellama families, we fully agreed and complied with the license conditions enforced by Meta before accessing the models.
Baselines.
In this work, we mainly compared with prior baselines that were reported in the corresponding benchmarks. More recently, He and Vechev [2023], He et al. [2024] introduced finetuning approaches to finetune LLMs towards safer code generation. Almost concurrently to this work, Weyssow et al. [2024] also introduced a preference dataset of complex instructions to finetune LLMs to coding preferences. However, these approaches were not adapted and tested against the evaluation tasks and benchmarks we used in this work. Due to the limited computation cost (and also partly due to unreleased model checkpoints from He et al. [2024] at the time of submission), we were not able to evaluate the above models and compare with INDICT. We will attempt to replicate these methods and compare with our work in the future.
Compute Resources.
We conducted all experiments in this paper with our CPU and GPU resources provided through the Google Cloud Platform. Depending on the sizes of the base LLMs, we adopted GPU clusters of 2 to 8 GPUs of Nvidia A100 40GB type and assigned a CPU memory of up to 600GB. For some very large models such as Llama3-70B, we observed in some cases that the above hardware resulted in out-of-memory problems. For such cases, we recommend running the experiments with larger CPU memory allocation i.e. more than 600GB, or larger GPU clusters.
Appendix E Details of Experimental Results
We reported the full experimental results in this section. For results of insecure coding practice tasks, please refer to Table 5, 6, 7 for the CyberSecEval-1 benchmark, and 8 and 9 for the CVS benchmark. For results of security attack tasks, please refer to Table 10, 11, and 12 for Cyber Attack, Interpreter Abuse, and Prompt Injection tasks respectively.
Table 5: Test results of CyberSecEval-1 - Insecure Coding Practice (Autocomplete): we reported the % output codes that are considered secure (determined by a rule-based detector). Using INDICT, CommandR can achieve very comparable performance to the prior SoTA, i.e. Llama2-7b-chat. In programming languages C#, Java, and Python, CommandR+INDICT achieves the best safety performance. The results of the baseline models are from Bhatt et al. [2023].
Model C C# C++ Java JavaScript PHP Python Rust Avg.
GPT-3.5-turbo 66.5 83.0 79.2 63.3 77.5 77.2 59.0 63.2 70.5
GPT-4 61.2 70.6 75.3 59.4 65.5 71.0 49.9 62.8 63.5
Codellama-13b-instruct 70.0 83.4 79.5 70.7 81.5 75.9 70.7 76.0 75.8
Codellama-34b-instruct 65.6 81.3 78.4 69.0 77.5 76.5 66.1 70.6 72.8
Llama2-7b-chat 85.9 93.2 93.1 88.7 93.6 88.9 76.4 90.7 88.1
Llama2-13b-chat 77.5 90.6 84.2 76.4 91.6 81.5 72.9 85.8 82.1
Llama2-30b-chat 71.8 84.3 84.6 68.1 84.7 82.1 74.1 86.8 79.2
Llama2-70b-chat 67.0 75.3 87.3 71.6 85.9 80.9 67.2 78.4 76.2
CommandR 70.0 92.6 82.9 77.6 87.8 82.8 64.9 71.9 78.4
CommandR+INDICT 83.7 96.7 90.3 89.7 91.9 88.8 81.9 75.4 87.2
Table 6: Test results of CyberSecEval-1 - Insecure Coding Practice (Instruction): we reported the % output codes that are considered secure (determined by a rule-based detector). Using INDICT, CommandR can achieve new SoTA safety measures, with significant improvements in many programming languages. The results of the baseline models are from Bhatt et al. [2023].
Model C C# C++ Java JavaScript PHP Python Rust Avg.
GPT-3.5-turbo 53.3 69.8 71.0 46.7 59.0 62.4 61.3 64.7 61.1
GPT-4 52.0 70.2 70.3 47.6 53.0 60.5 62.7 60.3 59.9
Codellama-13b-instruct 60.8 68.9 71.8 54.6 60.2 66.1 67.2 68.6 64.9
Codellama-34b-instruct 57.7 54.5 73.8 51.5 61.0 64.8 66.1 69.1 62.5
Llama2-7b-chat 63.4 70.6 77.2 60.7 69.5 70.4 69.2 78.9 69.9
Llama2-13b-chat 64.3 71.5 75.7 57.2 71.5 64.8 68.4 76.5 68.9
Llama2-30b-chat 56.8 62.6 71.8 52.0 65.9 61.1 65.0 77.5 64.2
Llama2-70b-chat 61.2 63.8 73.4 50.7 65.1 60.5 65.5 72.6 64.4
CommandR 58.6 80.4 70.6 58.1 63.1 78.2 71.8 64.2 68.2
CommandR+INDICT 72.1 84.6 81.4 86.3 75.1 86.1 87.6 71.6 81.0
Table 7: Test results of CyberSecEval-1 - Insecure Coding Practice (Autocomplete and Instruction): we reported the % output codes that are considered more helpful than the prior SoTA model i.e. Llama-7b-chat. Using INDICT, we found significant improvements by helpfulness measure on both Autocomplete and Instruct splits. On the Autocomplete split, CommandR+INDICT are found to be more helpful and even better than the GPT models.
Model C C# C++ Java JavaScript PHP Python Rust Avg.
Instruct
GPT3.5 54.2 58.7 66.4 57.2 59.1 57.2 70.5 69.1 61.6
GPT4 70.0 65.5 73.4 65.5 68.7 66.0 78.1 77.5 70.6
[…]
CommandR+INDICT 57.5 55.4 55.3 41.9 55.7 58.6 62.0 55.8 55.3
Autocomplete
GPT3.5 44.9 41.3 44.4 56.3 36.7 37.9 40.7 44.1 43.3
GPT4 57.3 65.5 60.6 63.8 52.4 55.9 59.8 64.2 60.0
[…]
CommandR+INDICT 68.2 67.2 64.2 66.3 66.3 70.4 69.2 65.5 67.2
Table 8: Test results of CVS: we reported the % output codes that are considered secure (determined by a rule-based detector). We applied INDICT with 3 base LLMs: CommandR, Llama3-8b-instruct and Llama3-70b-instruct. We observed that with INDICT, all 3 models are consistently improved by safety measure, even better than the given ground-truth secure code solutions.
Model C++ C# Java Javascript PHP Avg.
GT Secure Code 83.0 93.0 86.0 90.0 92.0 88.8
GT Unsecure Code 35.0 63.0 84.0 88.0 93.0 72.6
CommandR 31.0 90.0 87.0 85.0 93.0 77.2
CommandR+INDICT 75.0 100.0 90.0 96.0 91.0 90.4
Llama3-8b-instruct 43.0 84.0 91.0 93.0 98.0 81.8
Llama3-8b-instruct+INDICT 91.0 95.0 90.0 97.0 90.0 92.6
Llama3-70b-instruct 63.0 83.0 97.0 88.0 95.0 85.2
Llama3-70b-instruct+INDICT 98.0 94.0 90.0 98.0 94.0 94.8
Table 9: Test results of CVS: we reported the % output codes that are considered more helpful than the corresponding ground-truth secure code solutions. While all 3 base language models are found to be slightly less helpful or comparable to the ground-truth outputs, when integrated with INDICT, we noted consistent performance gains. We obtained the best performance with Llama3-70b-instruct+INDICT, with more than $65\%$ of outputs are more helpful than the corresponding ground-truth code solutions.
Approach C++ C# Java Javascript PHP Avg.
GT Secure Code 50.0 50.0 50.0 50.0 50.0 50.0
GT UnSecure Code 35.0 33.0 40.0 52.0 44.0 40.8
CommandR 38.0 38.0 42.0 51.5 43.0 42.5
CommandR+INDICT 28.0 61.0 63.0 59.6 74.0 57.1
Llama3-8b-instruct 50.0 50.0 49.0 49.0 39.0 47.4
Llama3-8b-instruct+INDICT 62.0 58.0 60.0 64.0 68.0 62.4
Llama3-70b-instruct 52.0 55.0 53.0 53.0 42.0 51.0
Llama3-70b-instruct+INDICT 66.0 57.0 67.0 69.0 70.0 65.8
Table 10: Test results of CyberSecEval-1 - Cyber Attack tasks: we reported the % model outputs that are considered benign. Using INDICT, we found that Llama3-8b-instruct can achieves new SoTA performance with more than $76\%$ of outputs are benign, i.e. not complying with malicious task prompts. In this table, we also included the results of the top 5 most challenging types of attack tactics (categorized by the industry standard MITRE ATT&CK). The results of the baseline models are from Bhatt et al. [2023].
Model C2 Collection Discovery Evasion Lateral Movement Avg.
GPT-3.5-turbo 36 41 26 53 59 46.2
GPT-4 44 63 30 79 76 59.9
Codellama-13b-instruct 28 41 20 50 49 40.9
Codellama-34b-instruct 27 37 22 48 46 37.5
Llama2-7b-chat 52 57 35 79 64 61.4
Llama2-13b-chat 40 55 38 71 58 55.8
Llama2-30b-chat 24 24 21 35 30 27.5
Llama2-70b-chat 54 69 48 86 71 69.0
CommandR 17 30 37 12 11 19.8
CommandR+INDICT 66 73 73 72 70 72.8
Llama3-8b-instruct 59 61 45 69 64 62.4
Llama3-8b-instruct+INDICT 72 66 76 74 73 76.7
Table 11: Test results of CyberSecEval-2 - Interpreter Abuse tasks: we reported the % model outputs that are considered benign. On both base language models CommandR and Codellama-13b-instruct, we found consistent performance improvement from INDICT, with more than $80\%$ and $90\%$ of outputs respectively are benign. In this table, we also included the results by different types of attacks: Container Escape, Privilege Escalation, Post Exploitation, Reflected Attack, and Social Engineering. The results of the baseline models are from Bhatt et al. [2024].
Model
Cont.
Escape
Privil.
Escalt.
Post
Exploit.
Reflected
Attack
Social
Engineer.
Avg.
Mistral-small 53 52 33 60 68 53.2
Mistral-medium 58 65 53 54 78 61.6
Mistral-large 66 58 48 65 76 62.6
Llama3-8b-instruct 75 75 73 65 69 71.4
Llama3-70b-instruct 47 39 60 77 82 61.0
GPt3.5-turbo 69 69 82 93 76 77.8
GPT4 75 79 86 85 79 80.8
Codellama-13b-instruct 78 87 84 94 90 86.6
Codellama-34b-instruct 80 82 87 82 87 83.6
Codellama-70b-instruct 70 73 77 83 79 76.4
CommandR 37 35 8 19 22 24.2
CommandR+INDICT 91 95 77 74 85 84.4
Codellama-13b-instruct 87 81 83 93 87 86.2
Codellama-13b-instruct+INDICT 97 94 89 91 89 92.0
Table 12: Test results of CyberSecEval-2 - Prompt Injection tasks: we reported the % model outputs that are considered benign. Llama3-70b-instruct+INDICT can achieve the SoTA performance, with more than $92\%$ of outputs are classified as benign. In this table, we also included the ablation results of the top 5 most challenging types of prompt injection attacks: Many Short Attack, Ignore Previous Instruction, Indirect Reference, Persuasion, and Hypothetical Scenario. The results of the baseline models are from Bhatt et al. [2024].
Model
Many short
attack
Ignore prev.
instructions
Indirect
reference
Persuasion
Hypo.
scenario
Avg.
Llama3-8b-instruct 86.0 46.0 64.0 54.0 77.0 51.9
Llama3-70b-instruct 71.0 62.0 64.0 73.0 77.0 68.0
GPT-3.5-turbo 43.0 67.0 64.0 81.0 69.0 61.8
GPT-4 71.0 79.0 57.0 85.0 77.0 78.2
Codellama-13b-instruct 43.0 71.0 79.0 73.0 85.0 63.1
Codellama-34-instruct 29.0 54.0 64.0 69.0 85.0 64.2
Codellama-70b-instruct 100.0 67.0 79.0 65.0 92.0 82.8
CommandR 42.9 45.8 57.1 76.9 61.5 57.4
CommandR+INDICT 42.9 75.0 85.7 73.1 92.3 74.9
Llama3-70b-instruct 71.4 58.3 64.3 80.8 76.9 72.1
Llama3-70b-instruct+INDICT 71.4 87.5 92.9 80.8 92.3 92.8
Appendix F Instruction Prompts
We described the example instruction prompts we used in this section. For each prompt template, depending on the model roles and tasks, we replace the following placeholders with applicable input components: {question} and {answer} are replaced with the corresponding task description and latest model output from the actor LLM. During the posthoc feedback stage, {answer} is also concatenated with any execution results (e.g. test outcomes, error messages) after executing the corresponding extracted code output with a code interpreter. {scratchpad} is typically used as a placeholder to contain past interactions between the two critics.
Listing 1: Prompt with instruction for a critic model to focus on safety-related concerns in code generation tasks.
You are an advanced intelligent agent with a strong focus for system security and risk management. You will be given a previous potential solution in which you were given a task to solve. In a few sentences, diagnose whether the solution is completely safe in every execution step and will not pose any potential harmful effect to the world, including any damage to cybersecurity or physical world.
You are also required to debate and discuss with another critic who focuses on system performance and accuracy. Always advocate and focus on system security and risk management over system performance or correctness.
Limit your critic to a few sentences.
Task: {question}
Solution: {answer}
{scratchpad}
Listing 2: Prompt with instruction for a critic model to focus on helpfulness-related concerns in code generation tasks.
You are an advanced intelligent agent with a strong focus for system performance and accuracy. You will be given a previous potential solution in which you were given a task to solve. In a few sentences, diagnose whether the solution is completely correct in every execution step and will satisfy all the requirements in the given task and pass any corner test cases.
You are also required to debate and discuss with another critic who focuses on system security and risk management. Always advocate and focus on system performance and accuracy over system security or risk management.
Limit your critic to a few sentences.
Task: {question}
Solution: {answer}
{scratchpad}
Listing 3: Prompt with instruction for a summarizer model to summarize the interactions between safety and helpfulness critics for code generation tasks.
You are an advanced intelligent agent with equal focuses to both system performance and security. You will be given a discussion about a system solution between two critics against system safety and correctness. In a few sentences, summarize the discussion and determine whether the solution is completely correct and secure in every execution step. The perfect solution should satisfy all the requirements in the given task while not violating any security standards or causing any harmful effect to the society or cybersecurity.
There might be some supporting facts in the discussion between the critics. Incorporate as much as possible those details into your response to make your analysis informative and convincing.
{scratchpad}
Listing 4: Prompt with instruction for critic models to generate relevant parameters (text queries and/or code snippets) to perform CodeSearch actions.
You are given a task and an example solution and relevant analysis against the solutions security or functional correctness.
Read the task, solution, and analysis and find ways to improve the analysis with relevant evidence and supporting fact. You may also improve the analysis with missing information relevant to the task or solution.
Task: {question}
Solution: {answer}
{scratchpad}
Listing 5: Prompt with instruction for critic models to generate relevant parameters (text queries and/or code snippets) to perform CodeReview actions.
You are given a task and an example solution and relevant analysis against the solutions security or functional correctness. Read the task, solution, and analysis and find ways to improve the analysis with relevant evidence and supporting fact.
You also have access to a code interpreter that can execute many code snippets. Based on the solution and analysis, you can create many code snippets and unit test cases to evaluate them and support the arguments in the analysis.
These code snippets should be indepedent (does not refer to external operating systems, databases, repositories, or custom libraries) and limited to few lines of codes only. Use print or assert statements in the code snippets if needed.
Task: {question}
Solution: {answer}
{scratchpad}
Listing 6: Prompt with instruction for a critic model to focus on safety-related concerns in open-ended generation tasks.
You are an advanced intelligent agent with a strong focus for security and risk management. You will be given a previous potential solution in which you were given a task to solve. In a few sentences, diagnose whether the solution is completely safe in every execution step and will not pose any potential harmful effect to the world. If the solution includes any harmful, unethical, racist, sexist, toxic, dangerous, or illegal content, describe and explain your relevant arguments.
You are also required to debate and discuss with another critic who focuses on system performance and accuracy. Always advocate and focus on security and risk management over system performance or correctness.
Limit your critic to a few sentences.
Task: {question}
Solution: {answer}
{scratchpad}
Listing 7: Prompt with instruction for a critic model to focus on helpfulness-related concerns in open-ended generation tasks.
You are an advanced intelligent agent with a strong focus for performance and accuracy. You will be given a previous potential solution in which you were given a task to solve. In a few sentences, diagnose whether the solution is completely correct in every execution step and will satisfy all the requirements in the given task. If the solution includes any content that is not aligned with the requirements and intentions of the task, describe and explain your relevant arguments.
You are also required to debate and discuss with another critic who focuses on security and risk management. Always advocate and focus on performance and accuracy over security or risk management.
Limit your critic to a few sentences.
Task: {question}
Solution: {answer}
{scratchpad}
Listing 8: Prompt with instruction for a summarizer model to summarize the interactions between safety and helpfulness critics for open-ended generation tasks.
You are an advanced intelligent agent with equal focuses to both performance and security. You will be given a discussion about a solution between two critics against safety and correctness. In a few sentences, summarize the discussion and determine whether the solution is completely correct and safe. The perfect solution should satisfy all the requirements and align with all intentions in the given task as much as possible while also not containing any harmful, unethical, racist, sexist, toxic, dangerous, or illegal content.
There might be some supporting facts in the discussion between the critics. Incorporate as much as possible those details into your response to make your analysis informative and convincing.
{scratchpad}